Given that
1 + 3 = 4
1 + 3 + 5 = 9
1 + 3 + 5 + 7 = 16
1 + 3 + 5 + 7 + 9 = 25
Determine a pattern and use it to find the sum of the first 300 odd numbers

Found 2 solutions by Alan3354, CharlesG2:
Answer by Alan3354(69443)

Given that
1 + 3 = 4
1 + 3 + 5 = 9
1 + 3 + 5 + 7 = 16
1 + 3 + 5 + 7 + 9 = 25
n = 5, sum = 5^2
Determine a pattern and use it to find the sum of the first 300 odd numbers
The sum is the square of the number of terms.
300^2 = 90000

Answer by CharlesG2(834)

Given that
1 + 3 = 4
1 + 3 + 5 = 9
1 + 3 + 5 + 7 = 16
1 + 3 + 5 + 7 + 9 = 25
Determine a pattern and use it to find the sum of the first 300 odd numbers
4 = 2^2
9 = 3^2
16 = 4^2
25 = 5^2
25 + 11 = 36 = 6^2
the pattern is n^2
300^2 = 90000